Shift Summary for Day Shift 1/16/00


Subject: Shift Summary for Day Shift 1/16/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:26:57 CST


        Shift Summary, Day shift, 16/Jan/2000

On Shift: Elizabeth T., Albert E.

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.5
Hours of E799 data (special runs not included): 6.5
Hours of special runs (laser scans, muon runs...): 0

Accounting for lost time:
Lost permit, quench at A11, changing tapes
========================================================================
Runs ended this shift: 15543
Run 15544 BEAM_799 in progress

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: BobH and Valeri to find permit.
========================================================================
Overview of shift:

Shift began with the permit still lost. BobH and Valeri went on
access and discovered that L2 signal was not reaching crate 31 due
to a bad connection at patch panel. Lost one hour to a quench at A11.
Very smooth data-taking for final six hours of shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15543 7:30 Access to find lost permit.
                8:20 Quench at A11.
                8:20 MCR resets NM2S1 -- current stops wandering.
                9:15 Beam returns.
                10:35 Clean up onsplit caches.
        374 kfrend
                14:00 Mount split tape UPM058.
                15544 Run ends, tapes full.
15544 15:25 Run begins.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:27:00 CST